# publications.bib

### v3

@@ 4.2. Perverse $t$-exactness
We recall that $\mathrm{Bun}_{G}^{b}\simeq[\ast/\mathcal{J}_{b}]$, where $\mathcal{J}_{b}:=\mathrm{Aut}(\mathcal{E}_{b})$ is the group diamond parameterizing automorphisms of the bundle $\mathcal{E}_{b}$ attached to $b\in B(G)$ on $X$. The diamond $\mathcal{J}_{b}$ has pure cohomological $\ell$-dimension over the base (in the sense of [FS, Definition IV.1.17]) equal to $\langle 2\rho_{G},\nu_{b}\rangle$, where $\nu_{b}$ is the slope homomorphism of $b$. Moreover, we have that $\mathrm{Bun}_{G}$ is cohomologically smooth of pure $\ell$-dimension equal to 00 over the base. This motivates the following definition.
@@
We define a perverse $t$-structure $(\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}),\phantom{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}))$ on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ such that $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ lies in $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (resp. $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$) if and only if $j_{b}^{*}(A)$ (resp. $j_{b}^{!}(A)$) sits in cohomological degrees $\leq\langle 2\rho_{G},\nu_{b}\rangle$ ($\geq\langle 2\rho_{G},\nu_{b}\rangle$).
For $\phi$ induced from a generic toral parameter, we write $(\phantom{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi},\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi})$ for the restriction of this $t$-structure to $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. Let $\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ denote the heart. We are almost ready to formulate our first big result. To do this, we need the following definition.
@@
We say that $\phi_{T}$ is weakly normalized regular if it is generic and if $\chi$ denotes the character attached to $\phi_{T}$ under local class field theory, we have, for all $w\in W_{G}$ non-trivial, that
(15) ${}\chi\otimes\delta_{B}^{1/2}\not\simeq(\chi\otimes\delta_{B}^{-1/2})^{w}$
holds and similarly for $\chi$ replaced by $\chi^{w_{0}}$. Similarly, we say $\phi_{T}$ is regular if for all $w\in W_{G}$ non-trivial we have that $\chi\not\simeq\chi^{w}$.
To motivate this, we recall that, since $\phi_{T}$ is weakly normalized regular, we have by [Ham2, Theorem 10.10] an object $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\in\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, which is a perverse filtered Hecke eigensheaf on $\mathrm{Bun}_{G}$, assuming 4.4 holds. Moreover, it is supported on the set of unramified elements and, for $b\in B(G)_{\mathrm{un}}$, its stalks are given by
$\mathrm{Red}_{b,\phi}^{\mathrm{tw}}:=\bigoplus_{w\in W_{b}}\rho_{b,w}[-\langle 2\rho_{G},\nu_{b}\rangle],$
where we recall that $\rho_{b,w}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$. In particular, by Proposition 4.5 it defines an object in the localized category $\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. To show the desired perverse $t$-exactness property, we would like to use the Hecke eigensheaf property of $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})$. Given a geometric dominant cocharacter $\mu$, we consider the highest weight tilting module $\mathcal{T}_{\mu}$ attached to $\mu$. We let
@@
be the Hecke operator attached to the representation $\mathcal{T}_{\mu}$, where $E_{\mu}$ denotes the reflex field of $\mu$. The sheaf $T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$ carries a filtration which, if it splits, guarantees an isomorphism $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\boxtimes r_{\mu}\circ\phi\simeq T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$, and we say that $\phi_{T}$ is $\mu$-regular ([Ham2, Definition 10.11]) if such a splitting exists. Here $r_{\mu}:\hat{G}\rightarrow\mathrm{GL}(\mathcal{T}_{\mu})$ is the map defined by the tilting module $\mathcal{T}_{\mu}$. The condition of being $\mu$-regular is guaranteed by the following stronger condition, using [Ham2, Theorem 1.17].
@@
are trivial for $\nu$,$\nu^{\prime}$ defining distinct $\Gamma$-orbits of weights in the highest weight tilting module $\mathcal{T}_{\mu}$.
@@
In particular, strong $\mu$-regularity implies $\mu$-regularity, and if we know strong $\mu$-regularity then it implies $\mu^{\prime}$-regularity for any $\mathcal{T}_{\mu^{\prime}}$ which occurs as a direct summand of the tensor product $\mathcal{T}_{\mu}^{\otimes n}$, by [Ham1, Proposition 10.12]. Also, as we will see, strong $\mu$-regularity is often implied by generic for some suitably chosen $\mu$.
More importantly, we can use this to deduce the following.
@@
For any $\phi$ induced from a generic $\phi_{T}$, assume, for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, the representations $\rho_{b,w}$ are semi-simple, and that Assumption 4.4 is true. Then we have a direct sum decomposition
@@
where $\mathrm{D}^{\mathrm{adm}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ denotes the subcategory of admissible complexes.
Moreover, for any $A\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$, we have that the !!! and * pushforwards agree with respect to the inclusion $j_{b}:\mathrm{Bun}_{G}^{b}\rightarrow\mathrm{Bun}_{G}$.
@@
The first part of the Proposition follows from the second part. To see this, we use the semi-orthogonal decomposition of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ into $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ via the excision spectral sequence. Using that the !!! and *-pushforwards agree for all objects $A\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}$, we see that the excision spectral sequence degenerates and the first part of the claim follows. To see the second part, we now use Proposition 4.5 to see that an object $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ can only be supported on the HN-strata $\mathrm{Bun}_{G}^{b}$ for $b\in B(G)_{\mathrm{un}}$, and that the restriction of $A$ to $\mathrm{Bun}_{G}^{b}$ has irreducible constituents valued in subquotients of the representations $\rho_{b,w}$ for $w\in W_{b}$ varying. For the representations $\rho_{b,w}$, we have the following.
@@
[Ham2, Proposition 11.13] For all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, the natural map
@@
is an isomorphism assuming $\phi_{T}$ is weakly normalized regular and assumption 4.4 is true.
So the !!! and * pushforwards agree on the $\rho_{b,w}$, and, since we are assuming the representations $\rho_{b,w}$ are semisimple, the claim follows for any constituent of $\rho_{b,w}$. This is enough to conclude the claim for any $A\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}^{b}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$ using the following claim.
@@
Assuming 4.4, for $\phi$ a generic parameter and any $A\in\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$ the cohomology of $A$ has finite length.
@@
By Assumption 4.4, we know that any irreducible constituent of the cohomology of $A$ is an irreducible constituent of $\rho_{b,w}$ for some $w\in W_{b}$. It follows by [Vig, p. II.5.13] that there are only finitely many possibilities for the irreducible constituents. Therefore, by choosing $K\subset G(\mathbb{Q}_{p})$ a sufficiently small open compact such that all these representations have an invariant vector, we deduce, since $A^{K}$ is a perfect complex by assumption, that $A$ must have finite length cohomology. ∎
@@
Let $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$ be a weakly normalized regular and regular. Suppose there exists a $\mu$ which is not fixed under any $w\in W_{G}$ and $\phi_{T}$ is $\mu$-regular. Then, $i_{B}^{G}(\chi)$ is irreducible.
@@
It follows, by [Ham2, Corollary 11.23] and the assumed $\mu$-regularity, that we have an isomorphism $i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w})=i_{B^{w}}^{G}(\chi)$ for all $w\in W_{G}$. Here $B^{w}$ is the conjugate of $B$ by $w$. We write $r_{B}^{G}$ for the normalized parabolic restriction functor. We recall that we are working with $\ell$-modular coefficients in possibly non-banal characteristic so $i_{B}^{G}(\chi)$ may have cuspidal constituents. In particular, we will need the following lemma.
@@
We apply second adjointness [DH1, Corollary 1.3] to the map
@@
$0\rightarrow\sigma\rightarrow i_{B}^{G}(\chi)\rightarrow\sigma^{\prime}\rightarrow 0$
Since parabolic restriction is exact (for example by using second adjointness), we get an exact sequence
$0\rightarrow r_{B}^{G}(\sigma)\rightarrow r_{B}^{G}i_{B}^{G}(\chi)\rightarrow r_{B}^{G}(\sigma^{\prime})\rightarrow 0$
@@
where the inequality follows from the geometric Lemma [Dat, Section 2.8] [^7] Note that this bound however fails without taking normalized restriction because of the aforementioned cuspidal constituents of $i_{B}^{G}(\chi)$ in non-banal characteristic (cf. [Dat, Page 48]).. By the previous lemma, we conclude that $\ell(r_{B}^{G}(\sigma))<W_{G}$. Now, since we know that $\sigma\subset i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w})$ for all $w\in W_{G}$, Frobenius reciprocity implies that we have non-zero maps $r_{B}^{G}(\sigma)\rightarrow\chi^{w}$ for all $w\in W_{G}$. This gives a contradiction by the regularity of $\chi$. ∎
@@
$T_{\mu}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})^{BW_{\mathbb{Q}_{p}}}$
for the Hecke operator attached to the highest weight tilting module $\mathcal{T}_{\mu}$ of highest weight $\mu$. Then the operator restricted to $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ is perverse $t$-exact if $\phi_{T}$ is weakly normalized regular, Assumption 4.4 is true, the $\rho_{b,w}$ are semi-simple for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, and $\phi_{T}$ is $\mu$-regular.
@@
Using Lemma 4.17, the commutation of Hecke operators with colimits, Proposition 4.5, Proposition 4.15, and semi-simplicity of the representations $\rho_{b,w}$, we can reduce to showing, for all $b\in B(G)_{\mathrm{un}}$, that if we consider the complex
$\mathrm{Red}^{\mathrm{tw}}_{b,\phi}:=\bigoplus_{w\in W_{b}}i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}[-\langle 2\rho_{G},\nu_{b}\rangle]\in\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
@@
$T_{\mu}(j_{b!}(\mathrm{Red}^{\mathrm{tw}}_{b,\phi}))\in\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
for the fixed $\mu$. However, $\mathrm{Red}^{\mathrm{tw}}_{b,\phi}$ are the stalks of the perverse filtered Hecke eigensheaf $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})$ and, since $\phi_{T}$ is $\mu$-regular by assumption, we have an isomorphism:
$T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))\simeq\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\boxtimes r_{\mu}\circ\phi\in\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}^{BW_{E_{\mu}}}.$

### v5

@@ 4.3. Perverse $t$-exactness
We recall that $\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$, where $\mathcal{J}_{b}:=\mathrm{Aut}(\mathcal{E}_{b})$ is the group diamond parameterizing automorphisms of the bundle $\mathcal{E}_{b}$ attached to $b\in B(G)$ on $X$, as in (9) though here we view it as a diamond over $\mathop{\rm Spd}(\overline{\mathbb{F}}_{p})$. The diamond $\mathcal{J}_{b}$ has pure cohomological $\ell$-dimension over the base (in the sense of [16, Definition IV.1.17]) equal to $\langle 2\rho_{G},\nu_{b}\rangle$, where $\nu_{b}$ is the slope homomorphism of $b$. Moreover, we have that $\mathrm{Bun}_{G}$ is cohomologically smooth of pure $\ell$-dimension equal to 0 over the base. This motivates the following definition.
@@
We define a perverse $t$-structure $(\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}),\phantom{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell}))$ on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ such that $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ lies in $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (resp. $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$) if and only if $j_{b}^{*}(A)$ (resp. $j_{b}^{!}(A)$) sits in cohomological degrees $\leq\langle 2\rho_{G},\nu_{b}\rangle$ (resp. $\geq\langle 2\rho_{G},\nu_{b}\rangle$) (See [10, Proposition 8.1.5] for a detailed existence proof).
For the rest of this section, we assume that $\ell$ is very decent in the sense of Definition 4.11 so that we may invoke the results of [24] and §4.1.
We let $\phi$ be a semisimple L-parameter, and we write $(\phantom{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi},\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi})$ for the restriction of this $t$-structure to the full subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. One can check that this gives rise to a well-defined $t$-structure on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ such that the localization map $(-)_{\phi}$ is $t$-exact. Indeed, this follows by applying Proposition A.3 to the two set of compact generators $\{j_{b!}(\mathrm{cInd}_{K}^{J_{b}(\mathbb{Q}_{p})}(\Lambda))\}$ and $\{j_{b\sharp}(\mathrm{cInd}_{K}^{J_{b}(\mathbb{Q}_{p})}(\Lambda))\}$ for $K\subset J_{b}(\mathbb{Q}_{p})$ a varying open pro-$p$ subgroup of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (See [16, Theorem V.4.1]) and using that localization is an exact functor. Here $j_{b\sharp}$ denotes the exceptional left adjoint to $j_{b}^{*}$ considered in [16, Proposition VII.7.2]. Moreover, one easily checks that this also gives rise to a $t$-structure after further intersecting with $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ the full subcategory of sheaves $A$ such that $j_{b}^{*}A$ satisfies the property that its invariants under any pro-$p$ group $K\subset J_{b}(\mathbb{Q}_{p})$ is a perfect complex, by an analogous argument where one uses that the category of perfect complexes of $\overline{\mathbb{F}}_{\ell}$-vector spaces is preserved under standard truncation (more generally this holds for coefficients in a regular Noetherian ring). We let $\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ and $\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ denote the hearts of these $t$-structures.
One of the technical conditions introduced in [24] is the following.
@@
We say that a toral parameter $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$ is generic regular if it is generic in the sense of Definition 1.1 and $\phi_{T}$ is also regular in the sense that, if $\chi$ denotes the character attached to $\phi_{T}$ under local class field theory, we have that $\chi$ is regular, i.e for all $w\in W_{G}$ non-trivial, we have that
(19) $\chi\not\simeq\chi^{w}.$
We let $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ be the L-parameter induced by a toral parameter $\phi_{T}$. To motivate the condition, we recall that, if $\phi_{T}$ is generic regular, we have by [24, Theorem 9.10] an object $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\in\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, which is a perverse filtered Hecke eigensheaf on $\mathrm{Bun}_{G}$, assuming 4.5 holds. Moreover, it is supported on the set of unramified elements and, for $b\in B(G)_{\mathrm{un}}$, its stalks are given by
$\mathrm{Red}_{b,\phi}^{\mathrm{tw}}:=\bigoplus_{w\in W_{b}}\rho_{b,w}^{\chi}[-\langle 2\rho_{G},\nu_{b}\rangle],$
where we recall that $\rho_{b,w}^{\chi}:=i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}$ and $\chi$ is the character attatched to $\phi_{T}$ via local class field theory. In particular, by compatability of the Fargues-Scholze correspondence with local Langlands for torii and parabolic induction [16, Theorem I.9.6 (i), (viii)], it defines an object in the localized category $\mathrm{Perv}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. To show the desired perverse $t$-exactness property, we would like to use the Hecke eigensheaf property of $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})$. Given a geometric dominant cocharacter $\mu$ of $G$, we consider the highest weight tilting module $\mathcal{T}_{\mu}$ attached to $\mu$, we let
@@
be the Hecke operator attached to the representation $\mathcal{T}_{\mu}$, where $E_{\mu}$ denotes the reflex field of $\mu$. The sheaf $T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$ carries a filtration which, if it splits, guarantees an isomorphism $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\boxtimes r_{\mu}\circ\phi\simeq T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$, and we say that $\phi_{T}$ is $\mu$-regular ([24, Definition 9.11]) if such a splitting exists. Here $r_{\mu}:\hat{G}\rightarrow\mathrm{GL}(\mathcal{T}_{\mu})$ is the map defined by the tilting module $\mathcal{T}_{\mu}$. The condition of being $\mu$-regular is guaranteed by the following stronger condition, using [24, Theorem 9.10].
@@
are trivial for $\nu$$\nu^{\prime}$ defining distinct $\Gamma$-orbits of weights in the highest weight tilting module $\mathcal{T}_{\mu}$.
@@
In particular, strong $\mu$-regularity implies $\mu$-regularity, and if we know strong $\mu$-regularity then it implies $\mu^{\prime}$-regularity for any $\mathcal{T}_{\mu^{\prime}}$ which occurs as a direct summand of the tensor product $\mathcal{T}_{\mu}^{\otimes n}$, by [27, Proposition 10.12], and this uses the second part of Definition 4.11 (2) in a key way. Also, as we will see, strong $\mu$-regularity is often implied by genericity for some suitably chosen $\mu$.
More importantly, this condition can be used to deduce the following semi-simplicity result for the inductions $\rho_{b,w}$.
@@
For any $\phi$ induced from a generic regular $\phi_{T}$ we assume, for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, that the representations $\rho_{b,w}$ are semi-simple, and that Assumption 4.5 is true. Then we have a direct sum decomposition
@@
where $\mathrm{D}^{\mathrm{adm}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\subset\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ denotes the subcategory of admissible complexes regarded as a full subcategory of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ by !!-pushforward.
Moreover, for any $A\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$, we have that the !! and * pushforwards agree with respect to the inclusion $j_{b}:\mathrm{Bun}_{G}^{b}\rightarrow\mathrm{Bun}_{G}$.
@@
The first part of the Proposition follows from the second part. To see this, we use the semi-orthogonal decomposition of $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ into $\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ via the excision triangles. Using that the !! and *-pushforwards agree for all objects $A\in\mathrm{D}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{F}}_{\ell})_{\phi}$, we see that the excision spectral sequence degenerates and the first part of the claim follows. To see the second part, we now use Proposition 4.6 to see that an object $A\in\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ can only be supported on the HN-strata $\mathrm{Bun}_{G}^{b}$ for $b\in B(G)_{\mathrm{un}}$, and that the restriction of $A$ to $\mathrm{Bun}_{G}^{b}$ has irreducible constituents valued in subquotients of the representations $\rho_{b,w}$ for $w\in W_{b}$ varying. For the representations $\rho_{b,w}$, we use the following claim of [24].
@@
[24, Proposition 10.12] For all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, the natural map
@@
is an isomorphism assuming $\phi_{T}$ is generic regular and Assumption 4.5 is true.
Therefore, we know the !! and * pushforwards agree on the $\rho_{b,w}$, and, since we are assuming the representations $\rho_{b,w}$ are semisimple, the claim follows for any constituent of $\rho_{b,w}$. This is enough to conclude the claim for any $A\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}^{b}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$ using the following claim.
@@
Assuming 4.5, for $\phi$ a generic parameter and any $A\in\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi}$ the cohomology of $A$ in the standard $t$-structure has finite length.
@@
By Proposition 4.6, we know that any irreducible subquotient of the cohomology of $A$ is an irreducible constituent of $\rho_{b,w}$ for some $w\in W_{b}$. It follows by [66, Section II.5.13] that there are only finitely many possibilities for the irreducible constituents of $\rho_{b,w}$. Therefore, by choosing $K\subset G(\mathbb{Q}_{p})$ a sufficiently small open compact such that all these representations have an invariant vector, we deduce, since $A^{K}$ is a perfect complex by assumption, that $A$ must have finite length cohomology. ∎
@@
Let $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$ be a generic regular paramater, and set $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$ to be the character attached to $\phi_{T}$ under class field theory. Suppose there exists a $\mu$ which is not fixed under any $w\in W_{G}$ and $\phi_{T}$ is $\mu$-regular. Then, $i_{B}^{G}(\chi)$ is irreducible.
@@
It follows, just as in the proof of [24, Corollary 10.22] and the assumed $\mu$-regularity, that we have an isomorphism $i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w})=i_{B^{w}}^{G}(\chi)$ for all $w\in W_{G}$. Here $B^{w}$ is the conjugate of $B$ by $w$. We write $r_{B}^{G}$ for the normalized parabolic restriction functor. We recall that we are working with $\ell$-modular coefficients in possibly non-banal characteristic so $i_{B}^{G}(\chi)$ may have cuspidal constituents. In particular, we will need the following Lemma.
@@
We apply second adjointness [12, Corollary 1.3] to the non-zero map
@@
$0\rightarrow\sigma\rightarrow i_{B}^{G}(\chi)\rightarrow\sigma^{\prime}\rightarrow 0.$
Since parabolic restriction is exact ([66, Section II.2.1]), we get an exact sequence
$0\rightarrow r_{B}^{G}(\sigma)\rightarrow r_{B}^{G}i_{B}^{G}(\chi)\rightarrow r_{B}^{G}(\sigma^{\prime})\rightarrow 0.$
@@
where the inequality follows from the geometric Lemma [11, Section 2.8] [^11] Note that this bound however fails without taking normalized restriction because of the aforementioned cuspidal constituents of $i_{B}^{G}(\chi)$ in non-banal characteristic (cf. [11, Page 48]).. By the previous Lemma, we conclude that $\ell(r_{B}^{G}(\sigma))<|W_{G}|$. Now, since we know that $\sigma\subset i_{B}^{G}(\chi)\simeq i_{B}^{G}(\chi^{w})$ for all $w\in W_{G}$, Frobenius reciprocity implies that we have non-zero maps $r_{B}^{G}(\sigma)\rightarrow\chi^{w}$ for all $w\in W_{G}$. This gives a contradiction by the regularity of $\chi$, since we have exhibited $|W_{G}|$ quotients of $r_{B}^{G}(\sigma)$ given by $\chi^{w}$ which are all distinct representations, by the regularity of the character $\chi$. ∎
@@
for the Hecke operator attached to the highest weight tilting module $\mathcal{T}_{\mu}$ of highest weight $\mu$, where $E_{\mu}$ denotes the reflex field of $E$. Then the operator restricted to $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$ is perverse $t$-exact if $\phi_{T}$ is generic regular, Assumption 4.5 is true, the $\rho_{b,w}$ are semi-simple for all $b\in B(G)_{\mathrm{un}}$ and $w\in W_{b}$, and $\phi_{T}$ is $\mu$-regular. Here $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ is the parameter induced by $\phi_{T}$.
@@
Using Lemma 4.20, the commutation of Hecke operators with colimits (which follows since it is a left adjoint (See the discussion preceding [16, Theorem I.7.2])), Proposition 4.6, Proposition 4.18, and semi-simplicity of the representations $\rho_{b,w}$, we can reduce to showing, for all $b\in B(G)_{\mathrm{un}}$, that if we consider the complex
$\mathrm{Red}^{\mathrm{tw}}_{b,\phi}:=\bigoplus_{w\in W_{b}}i_{B_{b}}^{J_{b}}(\chi^{w})\otimes\delta_{P_{b}}^{-1/2}[-\langle 2\rho_{G},\nu_{b}\rangle]\in\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
@@
$T_{\mu}(j_{b!}(\mathrm{Red}^{\mathrm{tw}}_{b,\phi}))\in\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$
for the fixed $\mu$. However, $\mathrm{Red}^{\mathrm{tw}}_{b,\phi}$ are the stalks of the perverse filtered Hecke eigensheaf $\mathrm{nEis}(\mathcal{S}_{\phi_{T}})$ given by [24, Theorem 9.10] and, since $\phi_{T}$ is $\mu$-regular then, by definition ([24, Definition 9.11]), we have an isomorphism:
$T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))\simeq\mathrm{nEis}(\mathcal{S}_{\phi_{T}})\boxtimes r_{\mu}\circ\phi\in\mathrm{Perv}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}^{BW_{E_{\mu}}}.$